(– 9) + (– 11) is greater than (– 9) – ( – 11). 
 (– 9) + (– 11) is greater than (– 9) – ( – 11). 
 

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

The given statement is False
(–9) + (–11) = - 19

But, (–9)–(–11) =–9 + 11 = 2
So, -19 < 2
Hence, (–9) + (–11) < (–9)–(–11)


Thanks
Shraddha Singh
2 years ago
Dear Student,
Let's check the following statement,
 
(-9)+(-11) = -9-11 = -20
 
Whereas, (-9)-(-11) = -9+11 = 2
 
So, (-9)+(-11) is smaller than (-9)-(-11).
Hence, the given statement is false.
 
Hope it helps!
Thank you